# FAM-HRI: Foundation-Model Assisted Multi-Modal Human-Robot Interaction Combining Gaze and Speech

### v1

@@ IV-A4 Human View Intention Alignment
After projecting gaze vectors onto a single image as $\mathbf{p}^{gaze}_{{t_{i+n}}}$ and generating the corresponding scene observation $\mathcal{Z}_{g}$, the next goal is to fuse gaze and language to infer the specific referred object $\overline{\mathcal{Z}_{g}}$ of the user. In our proposed approach, this intention fusion is determined by a weighted distance function by calculating the distance among each target object to all the gaze points, denoted as:
@@
Where $N=\left\lfloor\Delta t\times 20-1\right\rfloor$ and the weighting factor $e^{\alpha(n-N)}$ ensures that more recent gaze points contribute more to the final intention fusion.
@@ IV-C Planning Policy Generation
To translate this state space into an executable sequence of parameterized action primitives, we employ a GPT-based planning policy generator. We adapt GPT-4o with frozen checkpoints from 2024-08-06, using a temperature of 1 and input $X$ for policy generation. The model is prompted to generate semantic-level parameterized action primitives in JSON format, denoted as $\pi$.
To generate reasonable policy, we design structured prompts that enforce strict constraints on parameterized action primitives usage, the coordinate system, and the expected output format. By providing task descriptions, parameterized action primitive definitions, example tasks, and structured planning steps, we ensure that the LLM outputs consistent, interpretable, and physically feasible actions, reducing ambiguity and enhancing execution reliability. An example input and output of the planning policy generation is shown as follows:
@@
$\mathbf{O}_{1}$: $(object,\ position),(apple,\ table),(\Delta t_{apple},\ \Delta t_{there})$
$\overline{\mathcal{Z}_{r}}$: $(Position_{apple},\ Position_{there})$
$\mathcal{S}$: $please\ put\ the\ apple\ there\ on\ the\ table$
@@ V-A Perception and Manipulator Setup
Visual inputs for human view are collected through the Meta ARIA glasses [^3] https://www.projectARIA.com/, which also captures auditory signals through the built-in microphone as shown in Fig. 3. The robot view is collected through the Intel Realsense D435i RGBD camera. Our system has been tested with individuals in real-world environments, as illustrated in Fig. 4 (a-d). The scene includes a Franka Emika Panda 7DOF robot manipulator and various manipulation objects. An Intel D435i RGBD camera is positioned to monitor the workspace, ensuring clear visibility of both the environment and the robot. Multi-modal data processing is performed on a computer equipped with an RTX 4080 GPU, utilizing 6105 MB of GPU memory during interaction.
[…]
@@ V-B Description of Experimental Scenarios
To evaluate our system, we designed four experimental scenarios as shown in Fig. 4. Our system achieves an average LLM inference latency of ca. 1600 ms per command, with a worst-case latency of ca. 4200 ms. A full breakdown of inference latency under different conditions is provided in the supplementary materials.
@@
    Object Selection Among Similar Items: In the scenario 1 ($S_{1}$) shown in Fig. 4 (a), the test area contains nine visually similar pawns, each with a distance of 15cm between them, requiring the user to select a specific one.
@@
    As shown in Fig. 4 (b-d), Scenarios 2-4 ($S_{2}$-$S_{4}$) evaluate the execution of robotic tasks in increasingly complex real-world environments. The test area contains two similar cups, two similar exotic fruits (unknown name), and two similar plates. We categorize tasks into three levels of complexity: 1) Single-Step Action ($S_{2}$): Tasks requiring a single, direct action and where only one target object is referred to, such as picking up an object. 2) Multi-Step Action ($S_{3}$): Sequential actions requiring multiple steps to complete a task and where two target objects are referred to, such as put an object in a plate. 3) Causal Action ($S_{4}$): Tasks where an earlier action influences a later one, involving more than two target objects. For example, placing a fruit on a plate, then pouring liquid from a cup onto it without explicitly mentioning the plate’s position again.
Moreover, a user study was conducted to evaluate user satisfaction and experience. For the experiments and user study, we recruited 32 individuals, including 18 from Germany, and 14 from Singapore, all of whom are fluent in English.
@@ V-C Evaluation Metrics
To quantitatively assess system performance, we use two key metrics: success rate and interaction time. The success rate is defined as the proportion of correctly executed tasks ($N_{\text{correct}}$) to the total number of trials ($N_{\text{total}}$), given by $(N_{\text{correct}}/N_{\text{total}})*100\%$. Interaction time represents the efficiency of the system, measured as the time from when the user begins issuing a command to when the system completes interaction. This metric eliminates variations caused by different robot platforms, network situation, and computing configurations, ensuring a fair comparison across methods. The instruction of our proposed FAM-HRI and the survey can be found at Github.
@@ VI-A Baseline Selection
In our proposed FAM-HRI, we compare our proposed FAM-HRI with the various SOTA interaction methods utilizing different modalities. CaP [21] and ProgPrompt [22] are language-only interaction, making $S_{1}$ inefficient, as users would be required to specify the exact row and column of the pawns. GesSentence [4] utilizes gesture-only for interaction, HapGloves [34] employs a wearable tactile gloves for users to teleoperate the robot, while PhysicalDemo [35] leverages impedance-based feedback, where the user physically guides the robot through specific tasks. NVP-HRI [18] combines language and deictic posture for interaction, while GIRAF [19] fuses language with deictic gestures to achieve efficient object selection. These approaches are efficient in selecting similar objects, however, they rely on body movements and are less convenient for users with mobility or motor disabilities. FreeView [10] enables robot teleoperation through a virtual GUI controlled by the user’s gaze. Additionally since Gesture Sentence use pre-trained supervised networks for environment observation therefore cannot deal with unfamiliar objects in the scene. The results of baseline comparison, depicted in Tab. I.
Due to the lack of open-source code or similar hardware devices, we were unable to reproduce [10] and [34]. Instead, we directly report their success rate of task completion and interaction time from the original papers. However, since these methods used teleoperation, the comparison of success rate of task completion is not entirely fair and should be interpreted as a reference only.
@@
VI-B Object Selection Among Similar Items
In the $S_{1}$, as described in Sec. VB and illustrated in Fig. 4 (a), all 32 participants received verbal instructions on how to use FAM-HRI. To evaluate the our FAM-HRI distinguishing between multiple similar objects, each participant completed ten trials of pawn selection as instructed. The results, presented in Tab. I, show that our system achieved the second-highest success rate among all baseline methods in $S_{1}$, only lower than Physical Demo [35], which requires the user to physically guide the robotic arm near the target object. However, FAM-HRI significantly outperformed gesture-only and language-only methods, demonstrating its reliability in challenging and cluttered environments.
@@ VI-C Task Execution in Complex Environment
In $S_{2}-S_{4}$, described in Sec. VB and illustrated in Fig. 4 (b-d), the same 32 participants were asked to complete three levels of tasks using objects in the scene ($S_{2}$: Single-Step Action, $S_{3}$: Multi-Step Action, $S_{4}$: Causal Action). Each participant performed ten trials per task level.
@@ VI-D Comparison of Effectiveness
VI-E Qualitative Field Survey on User Experience
@@
To rigorously assess the proposed method, we conducted a user survey to evaluate interaction experience, including participants with mild upper-limb impairments. Participants responded to questions using a five-point Likert scale, assessing FAM-HRI and baseline methods in terms of object selection accuracy, device flexibility, approach modernity, ease of use, and interaction efficiency. As shown in Fig. 5, users rated FAM-HRI higher than all baselines across qualitative measures. Notably, it outperformed baselines in efficiency and accuracy, with all participants favoring FAM-HRI for its seamless integration of gaze and language signals. Additionally, most users found our system more modern and easier to operate, emphasizing its intuitive design. The hands-free interaction enabled by ARIA glasses significantly reduced physical effort, offering a more accessible and user-friendly solution for human-robot collaboration, particularly benefiting individuals with limited mobility.
@@
Figure 5: Results of User Experience Survey. The answer of the 32 participants are centered around zero for better comparison.
VII Conclusion and Future works
@@
However, FAM-HRI has several limitations. The computational latency introduced by LLMs inference and VLM may affect real-time responsiveness, making deployment on embedded systems or low-compute devices challenging. Additionally, speech input noise may introduce errors in object selection. A significant source of error arises from loose-fitting ARIA glasses, which can lead to misalignment in gaze estimation. Furthermore, our multi-view alignment relies on SuperGlue-based feature matching, but its robustness under object occlusion has not been extensively analyzed, potentially limiting performance in complex scenes.
Future work will focus on real-time optimization, exploring lighter models for edge deployment, and incorporating additional sensor modalities to enhance system robustness and adaptability in diverse HRI scenarios.

### v2

@@ IV-A4 Human View Intention Alignment
After projecting gaze vectors onto a single image as $\mathbf{p}^{gaze}_{{t_{i+n}}}$ and generating the corresponding scene observation $\mathcal{Z}_{g}$, the next goal is to fuse gaze and language to infer the specific referred object $\overline{\mathcal{Z}_{g}}$ of the user. In our proposed approach, this intention fusion is determined by a weighted distance function by calculating the distance between each target object to all the gaze points, denoted as:
@@
Where $N=\left\lfloor\Delta t\times 20-1\right\rfloor$ and the weight factor $e^{\alpha(n-N)}$ ensures that more recent gaze points contribute more to the final intention fusion.
Figure 4: Normalized Weights under Different $N$.
Figure 5: Comparison of Intention Alignment with and without the Weight Factor.
[…]
  1. 1.
    Ensuring Recent Gaze Points Have Higher Contribution: The exponential function $e^{\alpha(n-N)}$ weights recent gaze points higher while decaying older ones.
  2. 2.
    Avoiding Over-Reliance on a Single Gaze Point: The upper threshold of 0.65 prevents a single gaze point from dominating the selection.
@@ IV-B Multi-View Intention Alignment
Figure 6: Human view intention fusion and multi-view intention alignment. The user’s speech and gaze are aligned using a GPT-Agent, and the user intention is then processed by a multi-view alignment module to get the referred object within the robot manipulator view.
@@ IV-C Planning Policy Generation
To translate this state space into an executable sequence of parameterized action primitives, we employ a GPT-based planning policy generator. We adapt GPT-4o with frozen checkpoints from 2024-08-06, using a temperature of 1 and input $X$ for policy generation. The model is prompted to generate semantic-level parameterized action primitives in JSON format, denoted as $\pi$. In particular, our action primitives include both composite actions and atomic actions to flexibly support a wide range of task compositions:
  1. 1.
    Composite actions involve task-oriented behaviors commonly encountered in daily HRI scenarios:
    1. (a)
      Pick: grasp an object and lift it from the surface.
    2. (b)
      Put: place an object at a designated location.
    3. (c)
      Pour: tilt and empty the contents of a cup into a target container.
    4. (d)
      Swap: exchange the positions of two objects.
    5. (e)
      MoveTo: move the end-effector to a position.
  2. 2.
    Atomic actions offer fine-grained control over the robot’s motion and end-effector state:
    1. (a)
      MoveX, MoveY, MoveZ: move the end-effector along the x, y, or z axes.
    2. (b)
      OpenGripper: open the gripper.
    3. (c)
      CloseGripper: close the gripper.
    4. (d)
      Rotate: rotate the end-effector.
To generate a reasonable policy, we design structured prompts that enforce strict constraints on parameterized action primitives usage, the coordinate system, and the expected output format. By providing task descriptions, parameterized action primitive definitions, example tasks, and structured planning steps, we ensure that the LLM outputs consistent, interpretable, and physically feasible actions, reducing ambiguity and enhancing execution reliability. An example input and output of the planning policy generation is shown as follows:
@@
$\mathbf{O}_{1}$: $(object,\ position),(apple,\ table),(\Delta t_{apple},\ \Delta t_{here})$
$\overline{\mathcal{Z}_{r}}$: $(Position_{apple},\ Position_{here})$
$\mathcal{S}$: $please\ put\ the\ apple\ here\ on\ the\ table$
@@ V-A Perception and Manipulator Setup
Visual inputs for human view are collected through the Meta ARIA glasses [^3] https://www.projectARIA.com/, which also capture auditory signals through the built-in microphone. Examples of the outputs of the sensors used in the experiment and the reference frames are shown in Fig. 7. The robot view is collected through the Intel Realsense D435i RGBD camera. Our system has been tested with individuals in real-world environments, as illustrated in Fig. 8 (a-d). The scene includes a Franka Emika Panda 7DOF robot manipulator and various manipulation objects. An Intel D435i RGBD camera is positioned to monitor the workspace, ensuring clear visibility of both the environment and the robot. Multi-modal data processing is performed on a computer equipped with a RTX 4080 GPU, utilizing 6105 MB of GPU memory during interaction.
@@ V-B Description of Experimental Scenarios
To evaluate our system, we designed four experimental scenarios as shown in Fig. 8.
@@
    Object Selection Among Similar Items: In the scenario 1 ($S_{1}$) shown in Fig. 8 (a), the test area contains nine visually similar pawns, each with a distance of 15 cm between them, requiring the user to select one of the pawns designated by the experiment organizer by gazing at the bottom center of the pawns. The gaze error was reported in this scenario.
@@
    As shown in Fig. 8 (b-d), Scenarios 2-4 ($S_{2}$-$S_{4}$) evaluate the execution of robotic tasks in increasingly complex real-world environments. The test area contains two similar cups, two similar exotic fruits (unknown name), and two similar plates. We categorize tasks into three levels of complexity: 1) Single-Step Action ($S_{2}$): Tasks requiring a single, direct action and where only one target object is referred to, such as picking up an object. 2) Multi-Step Action ($S_{3}$): Sequential actions requiring multiple steps to complete a task and where two target objects are referred to, such as putting an object on a plate. 3) Causal Action ($S_{4}$): Tasks where an earlier action influences a later one, involving more than two target objects. For example, placing a fruit on a plate, then pouring liquid from a cup onto it without explicitly mentioning the plate’s position again.
Moreover, a user study was conducted to evaluate user satisfaction and experience. Fig. 9 shows the experimental setup. For the experiments and user study, a total of 32 participants were recruited, comprising 18 participants from Germany and 14 from Singapore. All participants are affiliated with local universities and fluent in English. Among the German participants, none have physical impairments; the group includes 4 women and 14 men, with one individual over the age of 60 and the rest under 50. Among the Singaporean participants, 6 have mild upper-limb impairments; the group consists of 6 women and 8 men, with two individuals over the age of 60 and the rest under 50. Each individual received only brief instructions and was asked to naturally move their eyes among the multiple referred objects during the experiment. Furthermore, we formulated hypotheses regarding user per- ceptions (H1, H2), and performed hypothesis testing to evaluate statistical significance.
Figure 9: Experimental Setup for user study. A disabled individual equipped with Aria glasses issues commands to the robot manipulator.
@@ V-C Evaluation Metrics
To quantitatively assess system performance, we use two key metrics: success rate and interaction time. The success rate is defined as the proportion of correctly executed tasks ($N_{\text{correct}}$) to the total number of trials ($N_{\text{total}}$), given by $(N_{\text{correct}}/N_{\text{total}})*100\%$. Interaction time represents the efficiency of the system, measured as the time from when the user begins issuing a command to when the system completes interaction. This metric eliminates variations caused by different robot platforms, network situations, and computing configurations, ensuring a fair comparison across methods.
In addition, we define task complexity as the sum of (1) the number of predefined actions in the action primitives and (2) the number of parameters, where parameters encompass both the number of referred objects and the associated spatial specifications such as positions and angles.
We have also included a supplementary video that presents experimental results, illustrating the system’s performance across various tasks and environments.
@@ VI-A Baseline Selection
In our proposed FAM-HRI, we compare our FAM-HRI with the various SOTA interaction methods utilizing different modalities. CaP [21] and ProgPrompt [22] are language-only interactions, making $S_{1}$ inefficient, as users would be required to specify the exact row and column of the pawns. GesSentence [4] utilizes gesture-only interaction, HapGloves [35] employs wearable tactile gloves for users to teleoperate the robot, while PhysicalDemo [36] leverages impedance-based feedback, where the user physically guides the robot through specific tasks. NVP-HRI [5] combines language and deictic posture for interaction, while GIRAF [19] fuses language with deictic gestures to achieve efficient object selection. These approaches are efficient in selecting similar objects. However, they rely on body movements and are less convenient for users with mobility or motor disabilities. FreeView [11] enables robot teleoperation through a virtual GUI controlled by the user’s gaze. Additionally, since Gesture Sentence use pre-trained supervised networks for environment observation, therefore cannot deal with unfamiliar objects in the scene. Comparison results are shown in Tab. I.
@@
Remark: Due to the lack of open-source code for [11]. We directly report their success rate of task completion and interaction time from the original papers.
“-” indicates that no relevant data are given explicitly or implicitly in the paper.
VI-B Object Selection Among Similar Items and Gaze Error
In the $S_{1}$, as described in Sec. V-B and illustrated in Fig. 8 (a), all 32 participants received verbal instructions on how to use FAM-HRI. To evaluate our FAM-HRI distinguishing between multiple similar objects, each participant completed ten trials of pawn selection as instructed. The results, presented in Tab. I, show that our system achieved the second-highest success rate among all baseline methods in $S_{1}$, only lower than Physical Demo [36], which requires the user to physically guide the robotic arm near the target object. However, FAM-HRI significantly outperformed gesture-only and language-only methods, demonstrating its reliability in challenging and cluttered environments. We obtained the gaze error by measuring the distance between the gaze point and the center of the bottom of the pawn. With each participant wearing ARIA glasses correctly, the gaze error is $1.58\pm 0.62$ cm.
@@ VI-C Task Execution in Complex Environment
In $S_{2}-S_{4}$, described in Sec. V-B and illustrated in Fig. 8 (b-d), the same 32 participants were asked to complete three levels of tasks using objects in the scene ($S_{2}$: Single-Step Action, $S_{3}$: Multi-Step Action, $S_{4}$: Causal Action). Each participant performed ten trials per task level.
@@ VI-D Comparison of Effectiveness
VI-E Evaluation of Tabletop Task
TABLE II: Detailed simulation tabletop manipulation success rate (%) and interaction time (s) across different task scenarios
Success Rate (%) Interaction time (s) Number of parameters Number of actions Complexity
With referred object
pick up the $<$object$>$ 100 1.3 1 1 2
grab the pieces ($S_{1}$) 96 1.8 1 1 2
put the $<$object$>$ on the $<$plate$>$ 100 2.2 2 3 5
put the $<$object$>$ on the $<$plate$>$ then
pour some thing from the $<$cup$>$ on it 94 5.9 4 6 10
put this $<$object$>$ $<$there$>$ (Position on table) 89 2.5 2 3 5
put the $<$object1$>$ and $<$object2$>$ on the $<$plate$>$ 92 3.8 3 6 9
put the $<$object1$>$ on the $<$plate1$>$ then
put the $<$object2$>$ on the $<$plate2$>$ 90 6.1 4 6 10
grab the $<$object$>$ and lift up for $<$distance$>$ then
turn it for $<$angle$>$ degrees 96 8.3 3 3 6
Without referred object
pick up this 100 1.1 1 1 2
grab this ($S_{1}$) 87 0.8 1 1 2
put this on that 98 1.7 2 3 5
put this on that then pour something from this on it 96 4.2 4 6 10
put this $<$there$>$ (Position on table) 93 1.9 2 3 5
put this and this on that 96 3.1 3 6 9
put this on this then put this on that 92 4.2 4 6 10
grab this and lift it up for $<$distance$>$ then
turn it for $<$angle$>$ degrees 95 7.9 3 3 6
Remark: The listed commands are indicative examples. Our system supports a wide range of natural language expressions with similar semantic intention, allowing flexible and intuitive user interaction beyond fixed templates.
Table II presents the success rates and average interaction times for various tabletop manipulation tasks. A key observation is that tasks executed without a referred object generally exhibit faster interaction times due to shorter command lengths. However, in Scenario $S_{1}$, the success rate for commands without explicit object references is lower. This is primarily because small, detailed parts of the chess pawns were sometimes misdetected by VLM. Consequently, during multi-view alignment, insufficient keypoints in these small detected parts were found, reducing the accuracy of object selection. Additionally, as the length and complexity of the user’s command increase, the reasoning process required by the two GPT-Agents becomes more involved, which in turn leads to a decrease in success rate due to higher inference uncertainty.
VI-F Qualitative Field Survey on User Experience
@@
To rigorously assess the proposed method, we conducted a user survey to evaluate interaction experience, including participants with mild upper-limb impairments. All participants were asked to complete a survey at the end of the experiment. Participants responded to questions using a five-point Likert scale, assessing FAM-HRI and baseline methods in terms of object selection accuracy, device flexibility, approach modernity, ease of use, and interaction efficiency. As shown in Fig. 10, users rated FAM-HRI higher than all baselines across qualitative measures. Notably, it outperformed baselines in efficiency and accuracy, with all participants favoring FAM-HRI for its seamless integration of gaze and language signals. Additionally, most users found our system more modern and easier to operate, emphasizing its intuitive design.
This user study investigated two main hypotheses:
  1. 1.
    H1: FAM-HRI is preferred by participants compared to all other baseline methods.
  2. 2.
    H2: FAM-HRI is quantitatively more effective and accurate compared to all other baseline methods.
H1: Participants rated FAM-HRI higher than all baseline methods as “light and flexible” ($p<0.001$), “modern” ($p<0.001$), and “simple to use” ($p<0.001$).
H2: Participants further rated the execution of their personal preferences, FAM-HRI, as “accurate” ($p<0.001$) and “efficient” ($p<0.001$).
The results show that our proposed approach significantly reduced physical effort, offering a more accessible and user-friendly solution for human-robot collaboration.
@@
Figure 10: Results of User Experience Survey. The answers of the 32 participants are centered around zero for better comparison. Users rated FAM-HRI higher than all baselines across qualitative measures.
VI-G Task Execution under varying noise levels
Figure 11: System robustness evaluation under varying background noise. The background noise was introduced using a laptop. The system consistently achieved 100% success rate up to 80 dB.
To evaluate the robustness of FAM-HRI in real-world conditions, we conducted experiments under varying ambient noise levels. As shown in Fig. 11(a), the background noise was introduced using a laptop placed near the user, playing a pre-recorded dialogue video. The content of the video simulated a real indoor conversation scenario, ensuring that the noise characteristics are similar to those encountered in household or public settings. A sound level meter was positioned next to the user to measure the real-time noise intensity. The noise intensity ranged from 20 dB to 90 dB, covering common noise conditions, such as conversation, traffic, vacuum cleaner, and hairdryer. According to the Level Comparison Chart  [^4] http://ehs.yale.edu/sites/default/files/files/decibel-level-chart.pdf, hearing loss may result from sustained exposure to 90dB or more.
The results, presented in Fig. 11(b), illustrate that our proposed FAM-HRI maintains a high success rate even under increasing noise levels. Here, the user needs to complete a pick-placing task. The system consistently achieved 100% success rate up to 70 dB, indicating strong resilience to moderate noise conditions. Even at 90+ dB, where typical speech recognition systems often fail, FAM-HRI still maintained an 85% success rate, demonstrating its robustness in handling speech input under challenging conditions.
The strong noise resistance is attributed to two aspects:
  1. 1.
    Hardware: Unlike traditional microphones that are directly exposed to ambient noise, the ARIA glasses’ microphones are embedded within the frame. This placement significantly reduces the impact of background noise, allowing the user’s voice to be clearly captured, even in high-noise environments.
  2. 2.
    Software: Our two LLM-Agent frameworks further enhance robustness against background noise. The GPT-Agent for human view command processing incrementally analyzes task-relevant verbs, nouns, and pronouns in the user’s command, filtering out irrelevant words and recognized background noise before passing a cleaned and structured instruction to the agent for policy generation.
VI-H Latency of LLMs Inference
Tab. III presents the inference latency of the two GPT-Agents under different experimental scenarios. The reported values include the average response time for human view command processing and planning policy generation, measured across all user trials.
The results show that command processing latency varies between 1353ms and 2202ms, with increased latency in more complex scenarios due to additional language reasoning. Similarly, the policy generation latency ranges from 1281ms to 1997ms, reflecting the increasing complexity of task planning as the number of required action primitives grows.
Despite these variations, our proposed FAM-HRI maintains a reasonable inference latency, demonstrating its feasibility for human-robot collaboration.
TABLE III: Inference latency under different conditions
Sceranio Latency of Human View Command Processing Latency of Policy Generation
$S_{1}$ $1353\pm 753$ms $1281\pm 692$ms
$S_{2}$ $1624\pm 660$ms $1535\pm 946$ms
$S_{3}$ $1855\pm 725$ms $1803\pm 891$ms
$S_{4}$ $2202\pm 936$ms $1997\pm 861$ms
Remark: The reported latency values correspond to inference using a full-scale LLM without any quantization [37], distillation [38], or retrieval-augmented generation (RAG) [39]. Incorporating such optimization strategies is expected to yield substantial reductions in latency.
VII Conclusion
@@
VIII Limitations and Future Works
FAM-HRI has several limitations. The computational latency introduced by LLMs inference and VLM may affect real-time responsiveness, making deployment on embedded systems or low-compute devices challenging. Future work will explore trade-offs between accuracy and speed, such as fine-tuning parameter-efficient large models using Low-Rank Adaptation (LoRA) [40] and implementing retrieval-augmented generation (RAG) [41] with smaller backbones to reduce latency.
A further significant source of error arises from loose-fitting ARIA glasses, which can lead to misalignment in gaze estimation. Cluttered or outdoor environments, where lighting variations, occlusions, and background motion introduce additional uncertainty. To address these limitations, future work will focus on developing a VLM that explicitly incorporates diverse environmental conditions during training [41], enabling the system to generalize more robustly across a wider range of real-world HRI scenarios.
Moreover, our current framework does not explicitly handle rare but possible situations in which a user’s speech command and gaze trajectory refer to different objects—e.g., the user looks at a cup on the left while verbally referring to a similar cup on the right. In our current method, the robot naturally resolves such ambiguities by selecting the object with the lowest weighted distance to the gaze trajectory. While these conflicting cases are unlikely in practical settings, we acknowledge their existence and plan to address them in future iterations. One direction involves integrating a dialogue module [42] that allows the robot to perform clarification actions, such as hovering above the candidate object to verify user intention before task execution. This would enable the system to robustly disambiguate multi-modal commands in more complex or ambiguous contexts.
Appendix A Failure Modes
Figure 12: Failure Modes.
As shown in Fig. 12, the failure cases of our system can be categorized into four key aspects: human view input, scene observation, multi-view alignment, and LLM reasoning.
  1. 1.
    Human View Input: For human view input, there are two sources of failure, speech recognition errors and inaccurate gaze estimation. The speech recognition error in our system primarily due to misinterpretation of similar-sounding words. Gaze estimation errors are largely influenced by improper wearing of the glasses. Specifically, when the nose bridge of the ARIA glasses does not correctly placed on the user’s nose, the accuracy of gaze estimation decrease significantly. However, to mitigate this, users were instructed to adjust their glasses properly before proceeding with the experiment, maximise ensuring that such errors do not propagate into the system.
  2. 2.
    Scene Observation: For scene observation, failures mainly come from more or less detected objects. In Scenario $S_{1}$, where users need to select a specific chess piece without explicitly stating its category, smaller parts of the chess pieces were sometimes incorrectly detected as separate objects. Additionally, speech recognition errors could result in the misclassification of the target object, causing the system to detect an incorrect category and subsequently select the wrong object.
  3. 3.
    Multi-View Alignment: For multi-view alignment, the main reason for failure is insufficient or unclear feature correspondence between human and robot views. Feature matching using superglue becomes unreliable in the presence of weak object textures, repetitive patterns, or partial occlusions, leading to incorrect correspondences.
  4. 4.
    LLM reasoning: For LLM Reasoning, the primary failure modes come from formatting inconsistencies and hallucinations in the generated output. Since FAM-HRI requires the LLM’s response to strictly follow a predefined prompt format, any deviation renders the output unusable by subsequent system modules. In rare cases, the LLM misinterprets the required action parameters, leading to incorrect parameter values or an incorrect ordering of action arguments. Additionally, logic errors happen in policy generation due to LLM hallucinations. For instance, in placement tasks, the generated policy may not open the gripper after reaching the target position, leading to task failure. However, such errors were observed infrequently in our experiments, as structured prompt design and output constraints significantly reduced the occurrence of these issues.
Appendix B Parameter Tuning and Reproducibility
TABLE IV: List of Parameters
Grounding Dino SAM2
Box Threshold 0.3
Text Threshold 0.3
Grounding-Model Tiny
SAM2 Model Large
SuperGlue
Max Keypoints 10000
Keypoints Threshold 1e-5
Match Threshold 1e-5
Resize No
Weights indoor
Our approach involves several parameters that are manually set or empirically tuned based on experimental observations.The Weight Factor in intention alignment was determined through extensive testing to balance recent gaze points’ influence while preventing a single point from dominating the selection process. While these parameters have been carefully adjusted to optimize system performance, we acknowledge that they are not mathematically proven to be optimal.
To ensure reproducibility, we provide explicit parameter settings for key components of our system in Tab. IV, including Grounding Dino for Scene Observation and SuperGlue for Multi-View Alignment.
Appendix C List of symbols and their meanings
To provide a comprehensive reference for the mathematical notations used throughout our paper, we include Tab. V in the supplementary materials. This table details all symbols, spaces, sets, and corresponding meanings, covering essential aspects of frame transformations, gaze reconstruction, scene observations, intention fusion, control system variables, and policy generation.
The notation system is structured to facilitate clarity in understanding multi-view geometric transformations, gaze-language fusion, and LLM-based policy generation. Specifically, it includes:
  1. 1.
    Frame representations for robot base, cameras, and gaze.
  2. 2.
    Scene Observations for both human and robot view, ensuring a unified representation for multi-view alignment.
  3. 3.
    Control system parameters, including action primitives, parameters, and planning policies.
This structured notation serves as a foundation for understanding FAM-HRI’s multi-modal interaction system, enabling reproducibility and further development within the research community.
TABLE V: List of symbols and their meanings
Symbol Space Meaning
()$\prescript{r}{}{()}$ Frame: Robot Base
()$\prescript{c}{}{()}$ Frame: Robot Camera
()$\prescript{gc}{}{()}$ Frame: Glasses Camera
()$\prescript{gp}{}{()}$ Frame: Glasses Pupil
$\mathcal{K}$ $\mathbb{R}^{3\times 3}$ Basic Geometry: Camera Intrinsic Matrix
$\mathbf{p}$ $\mathbb{R}^{2}$ Basic Geometry: 2D Position of a 3D coordinate
$\mathbf{P}$ $\mathbb{R}^{3}$ Basic Geometry: 3D coordinate
${}^{gc}T_{gp}$ $\mathbb{SE}(3)$ Transformation from Glasses Camera Frame to Glasses Pupil Frame
${}^{s}T^{head}_{t_{i}}$ $\mathbb{SE}(3)$ Transformation Representing Head Pose at Time $t_{i}$
$\beta$ $\mathbb{R}^{4}$ Environmental Bounding Box Representation $\left([x_{\min},y_{\min},x_{\max},y_{\max}]\right)$
$M$ $\mathbb{R}^{H\times W}$ Environmental Segmentation Mask (binary or probability map)
$\mathcal{S}$ $\Sigma^{*}$ Human Input Speech Sequence in Text
$\mathcal{G}$ $\mathbb{R}^{H\times W}$ Human input of Eye tracking camera
$\mathcal{U}$ $\mathbb{R}^{H\times W*(t_{n}-t_{1})}$ Human view: Image from User-Centric Camera on glasses
$\mathcal{Z}_{H}$ $\Sigma^{*}\times\mathbb{R}^{H_{1}\times W_{1}\times(t_{n}-t_{1})}\times\mathbb{R}^{H_{2}\times W_{2}\times(t_{n}-t_{1})}$ Human Input State Vector / Descriptor
${}^{gc_{t_{i}}}\mathbf{P}_{{t_{i}}}$ $\mathbb{R}^{3}$ Human view: 3D Gaze Vector at Time $t_{i}$ in Glasses Carmea Frame at Time $t_{i}$
${}^{gc_{t_{i}}}\mathbf{P}_{{t_{i+n}}}$ $\mathbb{R}^{3}$ Human view: 3D Gaze Vector at Time $t_{i+n}$ in Glasses Carmea Frame at Time $t_{i}$
$\mathcal{Z}_{g}$ Human view: Scene Observations Set $\left\{c_{target},\ \{\mathbf{p}_{g_{i}},\ \beta_{g_{i}},\ M_{g_{i}}\}_{i=1}^{N}\right\}$
$\mathbf{p}^{gaze}$ $\mathbb{R}^{2}$ Human View: 2D Projected Gaze Point on the Image Plane of $\mathcal{U}$
${}^{gc_{t_{i}}}\mathbf{p}^{gaze}_{{t_{i+n}}}$ $\mathbb{R}^{2}$ Human View: Projected Gaze Point at Time $t_{i+n}$ on Image at Time $t_{i}$
$\overline{\mathcal{Z}_{g}}$ $(\overline{\mathbf{p}_{g}},\ \overline{\beta_{g}},\ \overline{M_{g}})$ Human view: Referred Object
$\mathcal{C}$ $\mathbb{R}^{H\times W}$ Robot view: Image from Camera of Robot
$\mathcal{Z}_{r}$ Robot view: Scene Observations Set $\left\{c_{target},\ \{\mathbf{p}_{r_{i}},\ \beta_{r_{i}},\ M_{r_{i}}\}_{i=1}^{N}\right\}$
$\overline{\mathcal{Z}_{r}}$ $(\overline{\mathbf{p}_{r}},\ \overline{\beta_{r}},\ \overline{M_{r}})$ Robot view: Referred Object
$\mathbf{O}_{1}$$\left\{\mathcal{L},\ c_{target},\ \Delta t\right\}$ Control system: LLM Output for Gaze Time Period Prediction
$\mathcal{L}$ $\Sigma^{*}$ Control System: Target Property Descriptor (e.g., Object, Position)
$c_{target}$ $\Sigma^{*}$ Control System: Target Object Category from User Command
[…]
$X$$\left\{\overline{\mathcal{Z}_{r}},\ \mathbf{O}_{1},\ \mathcal{S}\right\}$ Control System: State Space
$\mathbf{O}_{2}$ $\pi$ Control System: LLM Output for Policy Generation
$\theta$ $\mathbb{R}^{n}$ Control System: Action Parameter Space (Rotation, Translation, position, etc.)
$a$ $\Sigma^{*}$ Control System: Action Space Primitives
$\mathcal{A}$$\left\{a(\theta)\right\}$ Control System: Set of Action Primitives
[…]
$\pi$ $\pi:X\mapsto\ \mathcal{A}\times\Theta$ Control System: Parameterized Planning Policy for Executing Actions